OpenStudy (anonymous):

why is it that you can factor a sum of two cubes, but not a sum of two square?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Here is an explanation to consider.\[\Large\bf\sf x^2+4=0\]Subtracting 4 from each side,\[\Large\bf\sf x^2=-4\]Taking the square root of each side,\[\Large\bf\sf x=\pm \sqrt{-4}\]Uh oh, we can't take the square root of a negative number! :( So no `real` solutions exist.

zepdrix (zepdrix):

If you include complex/imaginary solutions though, Taking the root gives us,\[\Large\bf\sf x=\pm 2i\] That's what happens when we factor the `sum of squares`. We get complex factors, not real ones.\[\Large\bf\sf x^2+4=(x+2i)(x-2i)\]

zepdrix (zepdrix):

If you haven't been introduced to imaginary numbers yet, then I apologize. This might seem a little confusing in that case hehe.

OpenStudy (anonymous):

lol thanks for your help this is one of the question on my quiz i have to write why ?

zepdrix (zepdrix):

hmm

OpenStudy (anonymous):



zepdrix (zepdrix):

I'll give you a little more info to think about. Word your answer the way that makes sense to you. It's your test! XD\[\Large\bf\sf x^3+8=0\]Again, solving for x,\[\Large\bf\sf x^3=-8\]We `CAN` take the cube root of a negative number. Doing so gives us,\[\Large\bf\sf x=-2\]